# Unmasking Puppeteers: Leveraging Biometric Leakage to Disarm Impersonation in AI-based Videoconferencing

### v1

@@ Abstract
AI-based talking-head videoconferencing systems slash bandwidth by sending a compact pose-expression latent and re-synthesizing RGB at the receiver—but this latent can be “puppeteered,” letting an attacker hijack a victim’s likeness in real time. Because every frame is synthetic, deepfake and synthetic video detectors fail outright. To address this security problem, we exploit a key observation: the pose expression latent inherently contain biometric information of the driving identity. Therefore, we introduce the first biometric leakage defense without ever looking at the reconstructed RGB video: a pose-conditioned, large-margin contrastive encoder that isolates persistent identity cues inside the transmitted latent while cancelling transient pose and expression. A simple cosine test on this disentangled embedding flags illicit identity swaps as the video is rendered. Our experiments on multiple talking-head generation models show that our method consistently outperforms existing puppeteering defenses, operates in real-time, and shows strong generalization to out-of-distribution scenarios.
@@ 3.2 Puppeteering Attacks
$\hat{V}_{t}^{k\rightarrow\ell}=g\bigl(f(V^{k}_{t}),R^{\ell}\bigr)=g\bigl(z^{k}_{t},R^{\ell}\bigr),$ (2)

### v2

@@ Abstract
AI-based talking-head videoconferencing systems reduce bandwidth by sending a compact pose-expression latent and re-synthesizing RGB at the receiver—but this latent can be “puppeteered,” letting an attacker hijack a victim’s likeness in real time. Because every frame is synthetic, deepfake and synthetic video detectors fail outright. To address this security problem, we exploit a key observation: the pose expression latent inherently contain biometric information of the driving identity. Therefore, we introduce the first biometric leakage defense without ever looking at the reconstructed RGB video: a pose-conditioned, large-margin contrastive encoder that isolates persistent identity cues inside the transmitted latent while cancelling transient pose and expression. A simple cosine test on this disentangled embedding flags illicit identity swaps as the video is rendered. Our experiments on multiple talking-head generation models show that our method consistently outperforms existing puppeteering defenses, operates in real-time, and shows strong generalization to out-of-distribution scenarios.Our code and trained models are available at https://github.com/MISLresearch/Unmasking-Puppeteers-Neurips25.
@@ 3.2 Puppeteering Attacks
$\hat{V}_{t}^{k\rightarrow\ell}=g\bigl(f(V^{\ell}_{t}),R^{k}\bigr)=g\bigl(z^{\ell}_{t},R^{k}\bigr),$ (2)
@@ 8 Conclusion
Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. 2320600. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.This research was supported in part by NVIDIA Corporation. We thank the NVIDIA Research team for their collaboration, resources, and valuable feedback throughout this project.